OpenStudy (anonymous):

perform the indicated operation. simplify. x^2-2x-3/x+1 divided by x^2 +x -12/x^2

OpenStudy (gw2011):

[(x^2-2x-3)/(x+1)]/(x^2+x-12)/x^2=[(x^2-2x-3)/(x+1)] [(x^2)/(x^2+x-12)] =(x-3)(x+1)(x^2)/(x+1)(x+4)(x-3)=x^2/(x+4)